OpenStudy (anonymous):

Solve the absolute value equation. Graph the solution. │3x + 1│ = 1

OpenStudy (anonymous):

Since it's absolute value, you'll need to solve if x was positive or if x was negative, so your equations are 3x+1=1 and 3x+1=-1.

OpenStudy (anonymous):

x=-2/3 or x=0?

OpenStudy (anonymous):

Since it's absoulte value, you usually always have 2 answers. So both!

OpenStudy (anonymous):

thats what i meant lol so they're right, ok.

OpenStudy (anonymous):

Yep!